1. A package is dropped from a helicopter flying at a height of 7.5 m while traveling 15 m/s. how far does the

package travel horizontally before it hits the ground below?

2. An orange rolls across with a speed of 0.5 m/s and falls off the edge. If it lands 3 m from the edge of the table,

from what height does it fall?


1
Expert's answer
2022-05-19T05:42:39-0400

1.

"L=v_0\sqrt{2H/g}=15*\sqrt{2*7.5/9.8}=19\:\rm m"

2.

"L=v_0\sqrt{2H/g}\\
H=g(L/v_0)^2/2=9.8*(3/0.5)^2/2=176\:\rm m"